(z+5)^2 = z^2 +25 ??????????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = (z+5)^2, which is similar to (a +b)^2identity, where (a+b)^2= a^2+b^2+2ab.
(z+5)^2= z^2+5^2+2×5×z = z^2+25+10z ≠z^2+25 = RHS
The correct statement is (z+5)^2= z^2+25+10z

Thanks